OpenStudy (anonymous):

Simplify the expression . 36 ^ -1/2

OpenStudy (anonymous):

1/6

OpenStudy (tyteen4a03):

@SGREBORN Please don't give out answers without explanations - this doesn't help the student learn and is in fact against OpenStudy's Code of Conduct. Remember that anything ^ -a = \(\frac{1}{anything^a}\). Deal with negative exponent first if possible.

OpenStudy (anonymous):

sorry the problem is 36^(1/2) and it needs to be simplified

OpenStudy (lacypennelll):

:|

OpenStudy (tyteen4a03):

Remember that anything ^ 1/2 = \(\sqrt(anything)\).